# Processing-in-memory for genomics workloads

### v1

@@ CONCLUSION
The initial promising results of BioPIM architectures, both within PnM and PuM frameworks, reveal that genome analysis can greatly benefit from the PIM approach and provide motivation for an active research field. Critical analysis of the results can give us valuable insights into the requirements and modifications required for current PIM platforms to support efficient genome analysis. We note that data-centric computing may also be achieved with near-storage-controller [18], in-flash [19], or near-router [20] computing, which we did not discuss in this paper.
While UPMEM’s current DPUs have proven the technology’s ability to speed up and reduce the cost and energy consumption of many applications compared to CPUs, GPUs, and FPGAs, they are limited by DRAM manufacturing constraints. These lightweight processors have limited computational power; for example, floating-point operations are emulated in software rather than implemented directly in hardware. To address the challenges of modern AI workloads, UPMEM is developing a dedicated PIM-AI chip [21]. This next-generation design uses a stacked configuration and integrates tensor and vector units to meet the needs of deep learning applications in all domains, including genomics.
@@
Community adoption of PIM technologies will require easy programmatic access to such devices for programmers who do not need to have an in-depth understanding of the underlying architectures. Application programming interfaces (APIs) and extensions to existing programming languages that hide hardware implementation from the programmer have made other accelerators widely used. For example, CUDA and OpenCL, in addition to APIs such as TensorFlow and PyTorch, are among the main driving forces for the immense popularity of GPGPU systems. Therefore, PIM technologies will also require such software layers for both general-purpose and domain-specific use cases. To this end, the BioPIM Consortium is currently developing a bioinformatics domain-specific API (BioPIM Library - BPL), in addition to a general-purpose API (SimplePIM), and an API for efficient computation of transcendental functions (TransPimLib).
Data production and associated analysis requirements are exponentially growing in all scientific domains. Beyond genomics, fields such as astronomy, high-energy particle physics, digital health, and image recognition are encountering a similar increase in data volume. Current methods to analyze such large amounts of data require large, power-hungry computing clusters and cloud infrastructure, where substantial amounts of energy and time are wasted in data movement. This increased energy demand, in turn, places additional strain on energy production capacities and potentially exacerbates challenges related to resource management. Therefore, there is a pressing need to co-develop efficient architectures and algorithms, especially for “big data” domains. Processing-in-memory technologies, while still being in their infancy, promise new opportunities to exploit to accelerate bioinformatics workloads while reducing energy requirements.

### v2

@@ CONCLUSION
The initial promising results of BioPIM architectures, both within PnM and PuM frameworks, reveal that genome analysis can greatly benefit from the PIM approach and provide motivation for an active research field. A critical analysis of the results can provide valuable insights into the requirements and modifications needed for current PIM platforms to support efficient genome analysis. We note that data-centric computing may also be achieved with near-storage-controller [18], in-flash [19], or near-router [20] computing, which we did not discuss in this paper.
While UPMEM’s current DPUs have proven the technology’s ability to speed up and reduce the cost and energy consumption of many applications compared to CPUs, GPUs, and FPGAs, they are limited by DRAM manufacturing constraints. These lightweight processors have limited computational power; for example, floating-point operations are emulated in software rather than implemented directly in hardware. To address the challenges of modern AI workloads, UPMEM proposed a dedicated PIM-AI chip [21]. This next-generation design uses a stacked architecture and integrates tensor and vector units to meet the needs of deep learning across all domains, including genomics.
@@
Community adoption of PIM technologies will require easy programmatic access to these devices for programmers who do not need an in-depth understanding of the underlying architectures. Application programming interfaces (APIs) and extensions to existing programming languages that hide hardware implementation from the programmer have made other accelerators widely used. For example, CUDA and OpenCL, along with APIs such as TensorFlow and PyTorch, are among the main drivers of the immense popularity of GPGPU systems. Therefore, PIM technologies will also require such software layers for both general-purpose and domain-specific use cases. To this end, the BioPIM Consortium is currently developing a bioinformatics domain-specific API (BioPIM Library - BPL), in addition to a general-purpose API (SimplePIM), and an API for efficient computation of transcendental functions (TransPimLib).
This manuscript demonstrates the acceleration of bioinformatic workloads on the UPMEM platform, validating real-world Processing-near-Memory (PnM) capabilities and our proposed PuM designs. While UPMEM represents a specific implementation, the insights gained here are transferable to emerging architectures such as HBM-PIM and CXL-PNM. These platforms share the critical ability to mitigate the data-movement bottleneck caused by high-latency external memory accesses. This architectural shift is essential because algorithmic evolution in genomics—ranging from Smith-Waterman alignment and variation calling to high-throughput base calling and metagenomic analysis—is creating stochastic memory access patterns that defeat standard caching hierarchies. Consequently, the gap left by traditional hardware limitations cannot be filled by simply increasing compute frequency; it requires adopting architectures that prioritize in-situ computation. To fully address these future use cases, research must focus on designing efficient communication frameworks for these commodity PIM devices, effectively decoupling genomic throughput from the limitations of memory bandwidth.
Data production and associated analysis requirements are growing exponentially across all scientific domains. Beyond genomics, fields such as astronomy, high-energy particle physics, digital health, and image recognition are encountering a similar increase in data volume. Current methods for analyzing such large amounts of data require large, power-hungry computing clusters and cloud infrastructure, where substantial energy and time are wasted on data movement. This increased energy demand, in turn, places additional strain on energy production capacities and potentially exacerbates challenges related to resource management. Therefore, there is a pressing need to co-develop efficient architectures and algorithms, especially for “big data” domains. Processing-in-memory technologies, while still being in their infancy, promise new opportunities to accelerate bioinformatics workloads while reducing energy requirements.